# Particle Semi-Implicit Variational Inference

### v1

@@ 5.3 Bayesian neural networks
Following prior works (e.g., Yu and Zhang (2023)), we compare our methods with other baselines on sampling the posterior of the Bayesian neural network for regression problems on a range of real-world datasets. We utilize the LSkip kernel $k_{\theta}(x|z)=\mathcal{N}(x;Wz+\mu_{\theta}(z),\sigma^{2}_{\theta}(z)I_{d_{x}})$. In Table 2, we show the root mean squared error on the test set. It can be seen that PVI performs well, or at least comparable, with other SIVI methods across all datasets. The details regarding the model and other parameters can be found Section H.3.
@@
Concrete (Yeh, 2007) $\bm{0.43}_{0.03}$ $0.50_{0.03}$ $0.50_{0.04}$ $0.92_{0.06}$
Protein (Rana, 2013) $\bm{0.90}_{0.15}$ $0.93_{0.08}$ $0.93_{0.07}$ $1.03_{0.06}$
Yacht (Gerritsma et al., 2013) $\bm{0.13}_{0.02}$ $0.17_{0.03}$ $0.18_{0.02}$ $0.98_{0.16}$
Table 2: Average Root Mean Square on the test set on various datasets. The results are averaged over $10$ independent trials. Here the $\pm$ denotes the average and we indicate in bold the lowest score.
@@ 6 Conclusion, Limitations, and Future Work
In this work, we frame SIVI as a minimization problem of $\mathcal{E}_{\lambda}$, and then, as a solution, we study its gradient flow. Through discretization, we propose a novel algorithm called Particle Variational Inference (PVI). Our experiments found that PVI can outperform current SIVI methods. At a marginal increase in computation cost (see Table 3) compared with prior methods, PVI can consistently perform better (or at least comparably in the worst cases considered) which we attribute to not imposing a particular form on the mixing distribution. This is a key advantage of PVI compared to prior methods: by not relying upon push-forward mixing distributions and instead using particles, the mixing distribution can express arbitrary distributions when the number of particles is sufficiently large. Furthermore, it is not necessary to tune the family of mixing distributions to obtain good results in particular problems. Theoretically, we study a related gradient flow of $\mathcal{E}_{\lambda}^{\gamma}$ and establish desirable properties such as the existence and uniqueness of solutions and propagation of chaos results. The main limitation of our work is that the theoretical results only apply to the case where $\gamma>0$; yet, our experiments were performed with $\gamma=0$ as this is when $\mathcal{E}_{\lambda}$ corresponds to the (regularized) evidence lower bound. Furthermore, we found that certain kernels were more amenable than others when it came to exploiting an expressive mixing distribution (e.g., the skip kernel). As such future work would include extending our analysis to cases where $\gamma=0$ and how to design effect kernels for PVI.
@@ Appendix A Related work
There are three algorithms for SIVI proposed: SVI Yin and Zhou [2018], UVI Titsias and Ruiz [2019], and SM Yu and Zhang [2023]. Each had their parameterization of SID (as discussed in Section 2), and their proposed optimization method. SVI relies on optimizing a bound of the ELBO which is asymptotically tight. UVI, like our approach, optimizes the ELBO by using gradients-based approaches. However, one of its terms is the score $\nabla_{x}\log q_{\theta,r}(x)$ which is intractable. The authors proposed using expensive MCMC chains to estimate it which, in contrast to PVI, this term is readily available to us. For SM, they propose to optimize the Fisher divergence, however, to deal with the intractabilities the resulting objective is a minimax optimization problem which is difficult to optimize compared to standard minimization problems.
@@
PVI utilizes the Euclidean–Wasserstein geometry. This geometry and associated gradient flows are initially explored in the context of (marginal) maximum likelihood estimation by Kuntz et al. [2023] and their convergence properties are investigated by Caprio et al. [2024]. In Lim et al. [2023], the authors investigated accelerated gradient variants of the aforementioned gradient flow in Euclidean–Wasserstein geometry. The Wasserstein geometry particularly for gradient flows on probability space has received much attention with many works exploring different functionals (for examples, see Arbel et al. [2019], Korba et al. [2021], Li et al. [2023]). In the context of variational inference Lambert et al. [2022] analyzed VI as a Bures–Wasserstein Gradient flow on the space of Gaussian measures.
@@
PVI is reminiscent of mixture distributions which is a consequence of the particle discretization. Mixture models have been studied in prior works as variational distributions[Graves, 2016, Morningstar et al., 2021]. In Graves [2016], the authors extended the parameterization trick to mixture distributions; and Morningstar et al. [2021] proposed to utilize mixture models as variational distributions in the framework of Kucukelbir et al. [2017]. Although similar, the mixing distribution assists the kernel in expressing complex properties of the true distribution at hand (see Section 5.1) which is an interpretation that mixture distribution lacks.
@@
There is an obvious similarity between SIVI and solving Fredholm equations of the first kind. There is considerable literature on solving such problems; see Crucinio et al. [2024], which is closest in spirit to the approach of the present paper, and references therein. In fact, writing $p(\cdot|y)=\int\tilde{k}(\cdot|z,\theta)r(z)\mathrm{d}z$. with $\tilde{k}(\cdot|z,\theta)\equiv k_{\theta}(\cdot|z)$ makes the connection more explicit: essentially, one seeks to solve a nonstandard Fredholm equation, with the LHS known only up to a normalizing constant, constraining the solution to be in $\mathcal{P}(\mathcal{Z})\times\{\delta_{\theta}:\theta\in\Theta\}$. While Crucinio et al. [2024] develop and analyse a simple Wasserstein gradient flow to address a regularised Fredholm equation, neither the method nor analysis can be applied to the SIVI problem because of this non-trivial constraint.
@@ Appendix B $\Gamma$-convergence
The following is one of many essentially equivalent definitions of $\Gamma$-convergence (see Dal Maso [2012], Braides [2002] for comprehensive summaries of $\Gamma$-convergence). We take as definition the following (see Dal Maso [2012, Proposition 8.1], Braides [2002, Definition 1.5]):
@@
Appendix C On Assumptions 3, 1 and 2
We shall show that the Gaussian kernel $k_{\theta}(x|z)=\mathcal{N}(x;\mu_{\theta}(z),\Sigma)$, i.e.,
@@ C.1 $k_{\theta}$ satisfies Assumption 2
Boundedness. First, we shall show with $k_{\theta}$ is bounded. Clearly, we have $k_{\theta}(x|z)\in\left[0,(2\pi)^{-d_{x}/2}\mathrm{det}(\Sigma)^{-0.5}\right]$ hence $|k_{\theta}|$ is bounded as a consequence of Assumption 4. Now to show that the gradient is bounded $||\nabla_{(\theta,x,z)}k_{\theta}(x|z)||$, we have the following
@@
Clearly, $\|\nabla_{(x,\mu,\sigma)}\mathcal{N}(x;\mu_{\theta}(z),\sigma^{2}I_{d_{x}})\|<\infty,$ from Assumption 4 and using the fact the gradient of a Gaussian density of given covariance w.r.t. $\mu$ is uniformly bounded.
Lipschitz. For $k_{\theta}$, one choice of coupling function and noise distribution is $\phi_{\theta}(z,\epsilon)=\Sigma^{\frac{1}{2}}\epsilon+\mu_{\theta}(z)$ and $p_{k}=\mathcal{N}(0,I_{d_{x}})$ where $\Sigma^{\frac{1}{2}}$ be the unique symmetric and positive definite matrix with $(\Sigma^{\frac{1}{2}})^{2}=\Sigma$ [Horn and Johnson, 2012, Theorem 7.2.6]; and the inverse map is $\phi_{\theta}^{-1}(z,x)=\Sigma^{-\frac{1}{2}}(x-\mu_{\theta}(z))$. Thus, from the change-of-variables formula, we have
@@
where $C$ is a constant and we use the following facts: $\|\tilde{\Sigma}^{-\frac{1}{2}}\|_{F}\leq|\mathrm{det}(\Sigma^{-1/2})|\|\Sigma^{-{1/2}}\|_{F}<\infty$ following from that fact $\Sigma^{-1/2}$ is positive definite; $p_{k}$ is a standard Gaussian with bounded derivatives hence is Lipschitz; and that the inverse map $\phi^{-1}$ is Lipschitz from Assumptions 4 and 5.
$\|\phi^{-1}_{\theta}(z,x)-\phi^{-1}_{\theta^{\prime}}(z^{\prime},x^{\prime})\|\leq\|\Sigma^{\frac{1}{2}}\|_{F}\|(x,\mu_{\theta}(z))-(x^{\prime},\mu_{\theta^{\prime}}(z^{\prime}))\|\leq C^{\prime}\|(x,\theta,z)-(x^{\prime},\theta^{\prime},z^{\prime})\|.$
@@ C.2 $k_{\theta}$ satisfies Assumption 3
The gradient is given by
$\nabla_{(\theta,z)}\phi_{\theta}(z,\epsilon):=\nabla_{(\theta,z)}\mu_{\theta}(z),$
@@
$p_{k}$ has finite second moments since it is a Gaussian.
@@ Appendix D Proofs in Section 2
where (a) follows the law of the unconscious statistician, and the last element-of follows from that fact that $k_{f;\phi,p_{k}}(\cdot|g(\epsilon)))=\phi(f\circ g(\epsilon),\cdot)_{\#}p_{k}\in\mathcal{K}_{\mathcal{F}\circ\mathcal{G};\phi,p_{k}}$. We can follow the argument above in reverse to obtain the reverse inclusion. Hence, we have obtained as desired.
As for $\mathcal{Q}_{\mathtt{YuZ}}=\mathcal{Q}_{\mathtt{YiZ}}$, follows in a similar manner, which we shall outline for completeness: let $q\in\mathcal{Q}(\mathcal{K}_{\mathcal{F};\phi,p_{k}},\mathcal{R}_{\mathcal{G},p_{r}})$, then
@@
One can follow in the reverse to obtain as desired. ∎
@@ Appendix E Proofs in Section 3
$\mathcal{E}(\theta,r)=\mathsf{KL}(q_{\theta,r}|p(\cdot|y))-\log p(y)\geq-\log p(y),$
@@
$\mathcal{E}(\theta,r)$ $=\int\underbrace{\mathds{1}_{[1,\infty)}\left(\frac{p(y,x)}{q_{\theta,r}(x)}\right)\log\left(\frac{q_{\theta,r}(x)}{p(y,x)}\right)}_{\leq 0}q_{\theta,r}(x)\,\mathrm{d}x$ (18)
$+\int\underbrace{\mathds{1}_{[0,1)}\left(\frac{p(y,x)}{q_{\theta,r}(x)}\right)\log\left(\frac{q_{\theta,r}(x)}{p(y,x)}\right)}_{\geq 0}q_{\theta,r}(x)\,\mathrm{d}x$ (19)
We shall focus on the RHS of (18). Applying Reverse Fatou’s Lemma, we obtain
$\limsup_{n\rightarrow\infty}-\int\mathds{1}_{[1,\infty)}\left(\frac{p(y,x)}{q_{\theta_{n},r_{n}}(x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(y,x)}\right)q_{\theta_{n},r_{n}}(x)\,\mathrm{d}x$
$\leq\int\limsup_{n\rightarrow\infty}\left(-\mathds{1}_{[1,\infty)}\left(\frac{p(y,x)}{q_{\theta_{n},r_{n}}(x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(y,x)}\right)q_{\theta_{n},r_{n}}(x)\right)\,\mathrm{d}x.$
@@
$\limsup_{n\rightarrow\infty}\mathds{1}_{[1,\infty)}\left(\frac{p(y,x)}{q_{\theta_{n},r_{n}}(x)}\right)$ $\leq\mathds{1}_{[1,\infty)}\left(\frac{p(y,x)}{q_{\theta,r}(x)}\right),$
$\lim_{n\rightarrow\infty}-\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(y,x)}\right)$ $=-\log\left(\frac{q_{\theta,r}(x)}{p(y,x)}\right),$
@@
where the first line is from u.s.c. of $\mathds{1}_{[1,\infty)}$; as for the second line, the continuity of $\log$; the final line follows the bounded kernel $k$ assumption and dominated convergence theorem.
@@
$\limsup_{n\rightarrow\infty}-\int\mathds{1}_{[1,\infty)}\left(\frac{p(y,x)}{q_{\theta_{n},r_{n}}(x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(y,x)}\right)q_{\theta_{n},r_{n}}(x)\,\mathrm{d}x$
$\leq$ $-\int\mathds{1}_{[1,\infty)}\left(\frac{p(y,x)}{q_{\theta,r}(x)}\right)\log\left(\frac{q_{\theta,r}(x)}{p(y,x)}\right)q_{\theta,r}(x)\,\mathrm{d}x,$
@@
$-\liminf_{n\rightarrow\infty}\int\mathds{1}_{[1,\infty)}\left(\frac{p(y,x)}{q_{\theta_{n},r_{n}}(x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(y,x)}\right)q_{\theta_{n},r_{n}}(x)\,\mathrm{d}x$
$\leq$ $-\int\mathds{1}_{[1,\infty)}\left(\frac{p(y,x)}{q_{\theta,r}(x)}\right)\log\left(\frac{q_{\theta,r}(x)}{p(y,x)}\right)q_{\theta,r}(x)\,\mathrm{d}x.$ (20)
Similarly, for the RHS of (19), using Fatou’s Lemma and using the l.s.c. of $\mathds{1}_{[0,1)}$, we obtain that
$\liminf_{n\rightarrow\infty}\int\mathds{1}_{[0,1)}\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(y,x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(y,x)}\right)q_{\theta_{n},r_{n}}(x)\,\mathrm{d}x$
$\geq$ $\int\mathds{1}_{[0,1)}\left(\frac{q_{\theta,r}(x)}{p(y,x)}\right)\log\left(\frac{q_{\theta,r}(x)}{p(y,x)}\right)q_{\theta,r}(x)\,\mathrm{d}x.$ (21)
@@
$\liminf_{n\rightarrow\infty}\mathcal{E}(\theta_{n},r_{n})=\liminf_{n\rightarrow\infty}\int\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(y,x)}\right)q_{\theta_{n},r_{n}}(x)\,\mathrm{d}x$
$\geq$ $\int\log\left(\frac{q_{\theta,r}(x)}{p(y,x)}\right)q_{\theta,r}(x)\,\mathrm{d}x\geq\mathcal{E}(\theta,r).$
@@
Consider the sequence $\Pi:=(\theta_{n},r_{n})_{n\in\mathbb{N}}$ where $\theta_{n}=\theta_{0}$ with $\|\theta_{0}\|<\infty$ $r_{n}=\delta_{n}$ with $k{{}_{\theta}}(x|z)=\mathcal{N}(x;\theta,I_{d_{x}})$ and $p(x|y)=\mathcal{N}(x;0,I_{d_{x}})$. Clearly, we have $q_{\theta,r}(x)=\mathcal{N}(x;\theta,I_{d_{x}})$ and so $\mathsf{KL}(q_{\theta,r}|p(\cdot|y))=\frac{1}{2}\|\theta\|^{2}.$ Hence, there is a $\beta<\infty$ such that
@@
Thus, we have shown that $\Pi\subset\{(\theta,r):\mathcal{E}(\theta,r)\leq\beta\}$. However, since the support of the elements of $\{r_{n}\in\mathcal{P}(\mathbb{R}^{d_{z}})\}_{n}$ eventually lies outside a ball of radius $R$ for any $R<\infty$ and hence of any compact set, $\Pi$ is not tight. Hence, Prokhorov’s theorem [Shiryaev, 1996, p. 318] tells us that as $\Pi$ is not tight it is not relatively compact. And, we conclude that as the level set is not relatively compact, the functional is not-coercive. ∎
@@
(Existence of a minimizer) The existence of a minimizer follows from coercivity and l.s.c. and applying Dal Maso [2012, Theorem 1.15]. ∎

### v3

@@ 5.4 Bayesian neural networks
Following prior works (e.g., Yu and Zhang, (2023)), we compare our methods with other baselines on sampling the posterior of the Bayesian neural network for regression problems on a range of real-world datasets. We utilize the LSkip kernel $k_{\theta}(x|z)=\mathcal{N}(x;Wz+f_{\theta}(z),\sigma^{2}_{\theta}(z)I_{d_{x}})$. In Table 2, we show the root mean squared error on the test set. It can be seen that PVI performs well, or at least comparable, with other SIVI methods across all datasets. The details regarding the model and other parameters can be found Section H.4.
@@
Concrete (Yeh,, 2007) $\bm{0.43}_{0.03}$ $0.50_{0.03}$ $0.50_{0.04}$ $0.92_{0.06}$
Protein (Rana,, 2013) $\bm{0.87}_{0.05}$ $0.92_{0.04}$ $0.92_{0.04}$ $1.02_{0.03}$
Yacht (Gerritsma et al.,, 2013) $\bm{0.13}_{0.02}$ $0.18_{0.02}$ $0.17_{0.02}$ $0.98_{0.16}$
Table 2: Root mean square error (lower is better) for Bayesian neural networks on the test set for various datasets. Here, we write the results in the form $\mu_{\sigma}$ where $\mu$ is the average RMS and $\sigma$ is its standard error computed over $10$ independent trials. We indicate in bold the lowest score.
@@ 6 Conclusion, Limitations, and Future Work
In this work, we frame SIVI as a minimization problem of $\mathcal{E}_{\lambda}$, and then, as a solution, we study its gradient flow. Through discretization, we propose a novel algorithm called Particle Variational Inference (PVI). Our experiments found that PVI can outperform current SIVI methods. At a marginal increase in computation cost (see Table 3) compared with prior methods, PVI can consistently perform better (or at least comparably in the worst cases considered) which we attribute to not imposing a particular form on the mixing distribution. This is a key advantage of PVI compared to prior methods: by not relying upon push-forward mixing distributions and instead using particles, the mixing distribution can express arbitrary distributions when the number of particles is sufficiently large. Furthermore, it is not necessary to tune the family of mixing distributions to obtain good results in particular problems. Theoretically, we study a related gradient flow of $\mathcal{E}_{\lambda}^{\gamma}$ and establish desirable properties such as the existence and uniqueness of solutions and propagation of chaos results.
The main limitation of our work is that the theoretical results only apply to the case where $\gamma>0$; yet, our experiments were performed with $\gamma=0$ as this is when $\mathcal{E}_{\lambda}$ corresponds to the (regularized) evidence lower bound. In future work, one can address these limitations by reducing this gap. Furthermore, we found that certain kernels were more amenable than others when exploiting an expressive mixing distribution (e.g., the skip kernel). The question of designing these kernels for PVI (or SIVI more generally) is important for future work.
Acknowledgments
JNL gratefully acknowledges the funding of the Feuer International Scholarship in Artificial Intelligence. AMJ acknowledges financial support from the United Kingdom Engineering and Physical Sciences Research Council (EPSRC; grants EP/R034710/1 and EP/T004134/1) and by United Kingdom Research and Innovation (UKRI) via grant EP/Y014650/1, as part of the ERC Synergy project OCEAN.
@@ Appendix A Related work
At the time of writing, there are three algorithms for SIVI proposed: SVI Yin and Zhou, (2018), UVI Titsias and Ruiz, (2019), SM Yu and Zhang, (2023). Concurrently, Cheng et al., (2024) extended the SM variant by solving the inner minimax objective and simplified the optimization problem. Each had their parameterization of SID (as discussed in Section 2), and their proposed optimization method. SVI relies on optimizing a bound of the ELBO which is asymptotically tight. UVI, like our approach, optimizes the ELBO by using gradients-based approaches. However, one of its terms is the score $\nabla_{x}\log q_{\theta,r}(x)$ which is intractable. The authors proposed using expensive MCMC chains to estimate it; in contrast to PVI, this term is readily available to us. For SM, they propose to optimize the Fisher divergence, however, to deal with the intractabilities the resulting objective is a minimax optimization problem which is difficult to optimize compared to standard minimization problems.
@@
PVI utilizes the Euclidean–Wasserstein geometry. This geometry and associated gradient flows are initially explored in the context of (marginal) maximum likelihood estimation by Kuntz et al., (2023) and their convergence properties are investigated by Caprio et al., (2024). In Lim et al., (2024), the authors investigated accelerated gradient variants of the aforementioned gradient flow in Euclidean–Wasserstein geometry. The Wasserstein geometry for gradient flows on probability space has received much attention with many works exploring different functionals (for examples, see Arbel et al., (2019); Korba et al., (2021); Li et al., (2023)). In the context of variational inference Lambert et al., (2022) analyzed VI as a Bures–Wasserstein Gradient flow on the space of Gaussian measures.
@@
PVI is reminiscent of mixture distributions which is a consequence of the particle discretization. Mixture models have been studied in prior works as variational distributions (Graves,, 2016; Morningstar et al.,, 2021). In Graves, (2016), the authors extended the parameterization trick to mixture distributions; and Morningstar et al., (2021) proposed to utilize mixture models as variational distributions in the framework of Kucukelbir et al., (2017). Although similar, the mixing distribution assists the kernel in expressing complex properties of the true distribution at hand (see Section 5.1) which is an interpretation that mixture distribution lacks.
@@
There is an obvious similarity between SIVI and solving Fredholm equations of the first kind. There is considerable literature on solving such problems; see Crucinio et al., (2024), which is closest in spirit to the approach of the present paper, and references therein. In fact, writing $p(\cdot|y)=\int\tilde{k}(\cdot|z,\theta)r(z)\mathrm{d}z$. with $\tilde{k}(\cdot|z,\theta)\equiv k_{\theta}(\cdot|z)$ makes the connection more explicit: essentially, one seeks to solve a nonstandard Fredholm equation, with the LHS known only up to a normalizing constant, constraining the solution to be in $\mathcal{P}(\mathcal{Z})\times\{\delta_{\theta}:\theta\in\Theta\}$. While Crucinio et al., (2024) develop and analyse a simple Wasserstein gradient flow to address a regularised Fredholm equation, neither the method nor analysis can be applied to the SIVI problem because of this non-trivial constraint. In Yan et al., (2024), the authors also solve a Fredholm-type equation but instead using the Wasserstein–Fisher–Rao geometry (Kondratyev et al.,, 2016; Gallouët and Monsaingeon,, 2017; Chizat et al.,, 2018; Liero et al.,, 2018).
@@ Appendix B $\Gamma$-convergence
The following is one of many essentially equivalent definitions of $\Gamma$-convergence (see Dal Maso, (2012); Braides, (2002) for comprehensive summaries of $\Gamma$-convergence). We take as definition the following (see Dal Maso, (2012, Proposition 8.1), Braides, (2002, Definition 1.5)):
@@
Appendix C On Assumptions 3 and 2
We consider the Gaussian kernel $k_{\theta}(x|z)=\mathcal{N}(x;\mu_{\theta}(z),\Sigma)$, i.e.,
@@ C.1 $k_{\theta}$ satisfies Assumption 2
In this section, we show that $k_{\theta}$ satisfies Assumption 2. We first show the boundedness property then the Lipschitz property.
Boundedness. First, we shall show that $k_{\theta}$ is bounded. Clearly, we have $k_{\theta}(x|z)\in\left[0,(2\pi)^{-d_{x}/2}\mathrm{det}(\Sigma)^{-0.5}\right]$ hence $|k_{\theta}|$ is bounded as a consequence of Assumption 4. Now to show that $||\nabla_{(\theta,x,z)}k_{\theta}(x|z)||$ is bounded, we have the following
@@
Hence, we have $\|\nabla_{(x,\mu,\sigma)}\mathcal{N}(x;\mu_{\theta}(z),\sigma^{2}I_{d_{x}})\|<\infty,$ from Assumption 4 and using the fact the gradient of a Gaussian density of given covariance w.r.t. $\mu$ is uniformly bounded. Thus, we have shown that $k_{\theta}$ satisfies the boundedness property in Assumption 2.
Lipschitz. For $k_{\theta}$, one choice of coupling function and noise distribution is $\phi_{\theta}(z,\epsilon)=\Sigma^{\frac{1}{2}}\epsilon+\mu_{\theta}(z)$ and $p_{k}=\mathcal{N}(0,I_{d_{x}})$ where $\Sigma^{\frac{1}{2}}$ be the unique symmetric and positive definite matrix with $(\Sigma^{\frac{1}{2}})^{2}=\Sigma$ (Horn and Johnson,, 2012, Theorem 7.2.6); and the inverse map is $\phi_{\theta}^{-1}(z,x)=\Sigma^{-\frac{1}{2}}(x-\mu_{\theta}(z))$. Thus, from the change-of-variables formula, we have
@@
where $C$ is a constant and we use the following facts: $\|\tilde{\Sigma}^{-\frac{1}{2}}\|_{F}\leq|\mathrm{det}(\Sigma^{-1/2})|\|\Sigma^{-{1/2}}\|_{F}<\infty$ following from the fact $\Sigma^{-1/2}$ is positive definite; $p_{k}$ is a standard Gaussian density function with Lipschitz gradients; and that the inverse map $\phi^{-1}$ is Lipschitz from Assumptions 4 and 5:
$\|\phi^{-1}_{\theta}(z,x)-\phi^{-1}_{\theta^{\prime}}(z^{\prime},x^{\prime})\|\leq\|\Sigma^{-\frac{1}{2}}\|_{F}\|(x,\mu_{\theta}(z))-(x^{\prime},\mu_{\theta^{\prime}}(z^{\prime}))\|\leq C^{\prime}\|(x,\theta,z)-(x^{\prime},\theta^{\prime},z^{\prime})\|.$
Hence, we have shown that $k_{\theta}$ satisfies the Lipschitz property of Assumption 2, and so Assumption 2 holds for $k_{\theta}$.
@@ C.2 $k_{\theta}$ satisfies Assumption 3
One can compute the gradient as
$\nabla_{(\theta,z)}\phi_{\theta}(z,\epsilon)=\nabla_{(\theta,z)}\mu_{\theta}(z),$
@@
$p_{k}$ has finite second moments since it is a Gaussian with positive definite covariance matrix.
@@ Appendix D Proofs in Section 2
where (a) follows the law of the unconscious statistician, and the last element-of follows from the fact that $k_{f;\phi,p_{k}}(\cdot|g(\epsilon)))=\phi(f\circ g(\epsilon),\cdot)_{\#}p_{k}\in\mathcal{K}_{\mathcal{F}\circ\mathcal{G};\phi,p_{k}}$. We can follow the argument above in reverse to obtain the reverse inclusion. Hence, we have obtained as desired.
That $\mathcal{Q}_{\mathtt{YuZ}}=\mathcal{Q}_{\mathtt{YiZ}}$, follows in a similar manner, which we shall outline for completeness: let $q\in\mathcal{Q}(\mathcal{K}_{\mathcal{F};\phi,p_{k}},\mathcal{R}_{\mathcal{G},p_{r}})$, then
@@
One can conclude by applying the same logic in the reverse direction. ∎
@@ Appendix E Proofs in Section 3
$\mathcal{E}(\theta,r)=\mathsf{KL}(q_{\theta,r},p(\cdot|y))-\log p(y)\geq-\log p(y),$
@@
$\mathcal{E}(\theta,r)$ $=\int\underbrace{\mathds{1}_{[1,\infty)}\left(\frac{p(x,y)}{q_{\theta,r}(x)}\right)\log\left(\frac{q_{\theta,r}(x)}{p(x,y)}\right)}_{\leq 0}q_{\theta,r}(x)\,\mathrm{d}x$ (20)
$+\int\underbrace{\mathds{1}_{[0,1)}\left(\frac{p(x,y)}{q_{\theta,r}(x)}\right)\log\left(\frac{q_{\theta,r}(x)}{p(x,y)}\right)}_{\geq 0}q_{\theta,r}(x)\,\mathrm{d}x$ (21)
We shall focus on the RHS of (20).
Note that we have the following bound
$\left|-\mathds{1}_{[1,\infty)}\left(\frac{p(x,y)}{q_{\theta_{n},r_{n}}(x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(x,y)}\right)q_{\theta_{n},r_{n}}(x)\right|$
$\leq$ $\max\left\{0,\log\left(\frac{p(x,y)}{q_{\theta_{n},r_{n}}(x)}\right)q_{\theta_{n},r_{n}}(x)\right\}\leq\max\{0,C\},$
where $C$ is some constant. The last inequality follows from the fact that the evidence is bounded from above and the kernel is bounded. We can apply Reverse Fatou’s Lemma to obtain
$\limsup_{n\rightarrow\infty}-\int\mathds{1}_{[1,\infty)}\left(\frac{p(x,y)}{q_{\theta_{n},r_{n}}(x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(x,y)}\right)q_{\theta_{n},r_{n}}(x)\,\mathrm{d}x$
$\leq\int\limsup_{n\rightarrow\infty}\left(-\mathds{1}_{[1,\infty)}\left(\frac{p(x,y)}{q_{\theta_{n},r_{n}}(x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(x,y)}\right)q_{\theta_{n},r_{n}}(x)\right)\,\mathrm{d}x.$
@@
$\limsup_{n\rightarrow\infty}\mathds{1}_{[1,\infty)}\left(\frac{p(x,y)}{q_{\theta_{n},r_{n}}(x)}\right)$ $\leq\mathds{1}_{[1,\infty)}\left(\frac{p(x,y)}{q_{\theta,r}(x)}\right),$
$\lim_{n\rightarrow\infty}-\log\left(\frac{p(x,y)}{q_{\theta_{n},r_{n}}(x)}\right)$ $=-\log\left(\frac{p(x,y)}{q_{\theta,r}(x)}\right),$
@@
where the first line is from u.s.c. of $\mathds{1}_{[1,\infty)}$; the second line from the continuity of $\log$; the final line follows from the bounded kernel $k$ assumption and dominated convergence theorem.
@@
$\limsup_{n\rightarrow\infty}-\int\mathds{1}_{[1,\infty)}\left(\frac{p(x,y)}{q_{\theta_{n},r_{n}}(x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(x,y)}\right)q_{\theta_{n},r_{n}}(x)\,\mathrm{d}x$
$\leq$ $-\int\mathds{1}_{[1,\infty)}\left(\frac{p(x,y)}{q_{\theta,r}(x)}\right)\log\left(\frac{q_{\theta,r}(x)}{p(x,y)}\right)q_{\theta,r}(x)\,\mathrm{d}x,$
@@
$-\liminf_{n\rightarrow\infty}\int\mathds{1}_{[1,\infty)}\left(\frac{p(x,y)}{q_{\theta_{n},r_{n}}(x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(x,y)}\right)q_{\theta_{n},r_{n}}(x)\,\mathrm{d}x$
$\leq$ $-\int\mathds{1}_{[1,\infty)}\left(\frac{p(x,y)}{q_{\theta,r}(x)}\right)\log\left(\frac{q_{\theta,r}(x)}{p(x,y)}\right)q_{\theta,r}(x)\,\mathrm{d}x.$ (22)
Similarly, for the RHS of (21), using Fatou’s Lemma (with varying measure and the set-wise convergence of $q_{\theta_{n},r_{n}}$) and using the l.s.c. of $\mathds{1}_{[0,1)}$, we obtain that
$\liminf_{n\rightarrow\infty}\int\mathds{1}_{[0,1)}\left(\frac{p(x,y)}{q_{\theta_{n},r_{n}}(x)}\right)\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(x,y)}\right)q_{\theta_{n},r_{n}}(x)\,\mathrm{d}x$
$\geq$ $\int\mathds{1}_{[0,1)}\left(\frac{p(x,y)}{q_{\theta,r}(x)}\right)\log\left(\frac{q_{\theta,r}(x)}{p(x,y)}\right)q_{\theta,r}(x)\,\mathrm{d}x.$ (23)
@@
$\liminf_{n\rightarrow\infty}\mathcal{E}(\theta_{n},r_{n})=\liminf_{n\rightarrow\infty}\int\log\left(\frac{q_{\theta_{n},r_{n}}(x)}{p(x,y)}\right)q_{\theta_{n},r_{n}}(x)\,\mathrm{d}x$
$\geq$ $\int\log\left(\frac{q_{\theta,r}(x)}{p(x,y)}\right)q_{\theta,r}(x)\,\mathrm{d}x\geq\mathcal{E}(\theta,r).$
@@
Consider the sequence $\Pi:=(\theta_{n},r_{n})_{n\in\mathbb{N}}$ where $\theta_{n}=\theta_{0}$; $\|\theta_{0}\|<\infty$; $r_{n}=\delta_{n}$; $k{{}_{\theta}}(x|z)=\mathcal{N}(x;\theta,I_{d_{x}})$; and $p(x|y)=\mathcal{N}(x;0,I_{d_{x}})$. Clearly, we have $q_{\theta,r}(x)=\mathcal{N}(x;\theta,I_{d_{x}})$ and so $\mathsf{KL}(q_{\theta,r},p(\cdot|y))=\frac{1}{2}\|\theta\|^{2}.$ Hence, there is a $\beta<\infty$ such that
@@
Thus, we have shown that $\Pi\subset\{(\theta,r):\mathcal{E}(\theta,r)\leq\beta\}$. However, since the support of the elements of $\{r_{n}\in\mathcal{P}(\mathbb{R}^{d_{z}})\}_{n}$ eventually lies outside a ball of radius $R$ for any $R<\infty$ and hence of any compact set, $\Pi$ is not tight. Hence, Prokhorov’s theorem (Shiryaev,, 1996, p. 318) tells us that, as $\Pi$ is not tight, it is not relatively compact. We conclude that, as the level set is not relatively compact, the functional is not-coercive. ∎
@@
(Existence of a minimizer) The existence of a minimizer follows from Dal Maso, (2012, Theorem 1.15) utilizing coercivity and l.s.c. of $\mathcal{E}_{\lambda}$. ∎